# Dimension-Free Estimates for Noncommutative Discrete Maximal Spherical Means

### v1

@@ Abstract.
In this paper, we establish dimension-free $L_{p}$-estimates for operator-valued maximal spherical means on cyclic groups ${\mathbb{Z}}_{m+1}^{d}$ for all $p>1$ and $m\geq 1$. The key ingredient is a noncommutative extension of the spectral technique developed by Nevo and Stein. As an application, we obtain a noncommutative spherical maximal inequality for automorphism actions of von Neumann algebras.
@@ 1. Introduction
The purpose of this paper is to establish the dimension-free $L_{p}$-bound (1) to the operator-valued setting, which then implies the corresponding noncommutative maximal inequalities. Let $\mathcal{M}$ be a von Neumann algebra equipped with a normal semifinite faithful trace $\tau$. The noncommutative $L_{p}$-spaces associated to the pair $({\mathcal{M}},\tau)$ is denoted by $L_{p}({\mathcal{M}})$. For each integer $k$, define the convolution operators
@@
for operator valued functions $f:{\mathbb{Z}}^{d}_{m+1}\rightarrow\mathcal{M}$. Let $\mathcal{N}=L_{\infty}({\mathbb{Z}}^{d}_{m+1},{\mathcal{M}})\cong L_{\infty}({\mathbb{Z}}^{d}_{m+1})\bar{\otimes}{\mathcal{M}}$ be the von Neumann algebra tensor product equipped with the tensor trace $\varphi=\mu\otimes\tau$, where $\mu$ is the Haar measure on ${\mathbb{Z}}^{d}_{m+1}$. Note that for $0<p<\infty$, $L_{p}(\mathcal{N})$ can be identified as the Bochner $L_{p}$-space $L_{p}({\mathbb{Z}}^{d}_{m+1};L_{p}({\mathcal{M}}))$. The following is our first main result.
@@
Let $(T_{k})_{1\leq k\leq d}$ be the convolution operators defined as (2). Then for all $p>1$ and for all operator-valued functions $f\in L_{p}(\mathcal{N})\cong L_{p}({\mathbb{Z}}^{d}_{m+1};L_{p}({\mathcal{M}}))$
@@
As noted in [14, Remark 1.6], the region $p>1$ is optimal because the optimal constant in the weak $(1,1)$ bound grows at least as fast as $\sqrt{d}$:
@@
An immediate difficulty in proving Theorem 1.1 is the fact that the classical maximal function $Mf=\sup_{1\leq k\leq d}|T_{k}f|$ does not make sense in general when $(T_{k}f)$ is a sequence of operators (see [17]). Indeed, the right formulation of noncommutative $L_{p}$-maximal inequalities was achieved via the introduction of vector-valued noncommutative $L_{p}$-spaces $L_{p}({\mathcal{M}};\ell_{\infty})$, which play the roles of the $L_{p}$-norm of maximal function $Mf$ in the noncommutative setting [16]. We refer to the next section for precise definitions.
As an application, we establish the noncommutative spherical maximal inequality for automorphism actions of ${\mathbb{Z}}_{m+1}^{d}$ on von Neumann algebras. Recall that for a von Neumann algebra $\mathcal{M}$}, we say that $\alpha:{\mathbb{Z}}_{m+1}^{d}\rightarrow\text{Aut}(\mathcal{M})$ is an action if for each $u$, $\alpha_{u}:\mathcal{M}\to\mathcal{M}$ is a *-preserving automorphism; and for all $s,u\in{\mathbb{Z}}_{m+1}^{d}$, $\alpha_{s}\circ\alpha_{u}=\alpha_{s+u}$. If in additional, $\tau\circ\alpha_{u}=\tau$ for all $u\in{\mathbb{Z}}_{m+1}^{d}$, we say $\alpha$ is a $\tau$-preserving action or an action of $\tau$-preserving automorphisms, denoted as $\alpha\curvearrowright(\mathcal{M},\tau)$. For each $k$, define
@@
Theorem 1.1, together with the noncommutative Calderón transference principle developed in [11], implies the following result. Readers may refer to Section 2 for the definition of $\ell^{c}_{\infty}$-valued $L_{p}(\mathcal{M})$ space $L_{p}({\mathcal{M}};\ell^{c}_{\infty})$.
@@
Let $\alpha$ be an $\tau$-preserving action of ${\mathbb{Z}}_{m+1}^{d}$ on a semi-finite von Neumann algebra $({\mathcal{M}},\tau)$. Let $(M_{k})_{1\leq k\leq d}$ be the defined as (3). Then for $p>1$,
@@
$\|(M_{k}x)_{1\leq k\leq d}\|_{L_{p}({\mathcal{M}};\ell^{c}_{\infty})}\leq\sqrt{C_{p,m}}\|x\|_{L_{p}({\mathcal{M}})},$
@@
Theorem 1.3 can be rephrased as a maximal inequality for multiple ergodic averages. Let $(L_{i})_{1\leq i\leq d}$ be a commuting family (i.e. $L_{i}L_{j}=L_{j}L_{i}$) of $\tau$-preserving automorphisms on $\mathcal{M}$} such that $L_{i}^{m+1}=id$. For each positive integer $k$, define
@@
We conclude this introduction by outlining the structure of the paper. Section 2 reviews some basic background on noncommutative $L_{p}$-spaces. The proof of Theorem 1.1 are presented in Sections 3 and 4. We prove Theorem 1.3 in Section 5.
@@ 2. Preliminaries
Throughout this paper, we denote by $C$ a generic positive constant whose value may change between occurrences. The notation $X\lesssim Y$ indicates $X\leq CY$ for some constant $C>0$ independent of essential parameters, while $X\thickapprox Y$ signifies that both $X\lesssim Y$ and $Y\lesssim X$ hold simultaneously.
@@ 2.1. Noncommutative $L_{p}$-spaces
Let $\mathcal{M}$} be a von Neumann algebra equipped with a normal semi-finite faithful trace $\tau:{\mathcal{M}}_{+}\rightarrow[0,\infty]$. Let $\mathcal{S_{{\mathcal{M}}+}}$ be the set of all positive elements $x\in{\mathcal{M}}$ such that $\tau(\mathrm{supp}(x))<\infty$, where $\mathrm{supp}(x)$ denotes the support projection of $x$ and $\mathcal{S}_{{\mathcal{M}}}$ be the linear span of $\mathcal{S_{{\mathcal{M}}+}}$. Then $\mathcal{S}_{{\mathcal{M}}}$ forms a $w^{*}$-dense $\ast$-subalgebra of $\mathcal{M}$}. For $1\leq p<\infty$, the noncommutative $L_{p}$-space $L_{p}({\mathcal{M}},\tau)$ (in short, $L_{p}(\mathcal{M})$) is the completion of $\mathcal{S}_{{\mathcal{M}}}$ under the norm,
@@ 2.2. Vector-valued noncommutative $L_{p}$-spaces
    $L_{p}({\mathcal{M}};\ell_{1})=\big(L_{p_{0}}({\mathcal{M}};\ell_{1}),L_{p_{1}}({\mathcal{M}};\ell_{1})\big)_{\theta}\ ,\ \text{isometrically}$
@@
Finally, we introduce the $\ell_{\infty}$-valued noncommutative weak $L_{p}$-spaces $\Lambda_{p,\infty}({\mathcal{M}};\ell_{\infty})$ (see e.g. [10]). Given a sequence $(x_{n})_{n}$ in $L_{p}(\mathcal{M})$ with $1\leq p<\infty$, we define
$\|(x_{n})_{n}\|_{\Lambda_{p,\infty}({\mathcal{M}};\ell_{\infty})}=\sup_{\lambda>0}\ \lambda\inf\Big\{\big(\tau(e^{\perp})\big)^{\frac{1}{p}}:e\in\mathcal{P}({\mathcal{M}})\text{ s.t. }\|ex_{k}e\|_{\infty}\leq\lambda\ \mbox{for\ all}\ k\Big\},$
where $\mathcal{P}({\mathcal{M}})$ is the set of all projections in $\mathcal{M}$}. The space $\Lambda_{p,\infty}({\mathcal{M}};\ell_{\infty})$ consists of all sequences $x=(x_{n})_{n}$ for which this quasi-norm is finite.
@@ 2.3. Noise operators
    $\|(H_{T}f)_{T>0}\|_{\Lambda_{1,\infty}(\mathcal{N};\ell_{\infty})}\leq C_{1,m}\|f\|_{L_{1}(\mathcal{N})};$
@@ 3. The smooth spherical maximal operators
    $\|(T_{K}^{M}f)_{0\leq K\leq\frac{md}{m+1}}\|_{\Lambda_{1,\infty}(\mathcal{N};\ell_{\infty})}\leq C_{1,m}\|f\|_{L_{1}(\mathcal{N})}\ ,$
    $\|(T_{K}^{D}f)_{0\leq K\leq\frac{d}{m+1}}\|_{\Lambda_{1,\infty}(\mathcal{N};\ell_{\infty})}\leq C_{1,m}\|f\|_{L_{1}(\mathcal{N})}\ ;$
@@
    $\|(J_{P}f)_{0<P\leq\frac{m}{m+1}}\|_{\Lambda_{1,\infty}(\mathcal{N};\ell_{\infty})}\leq C_{1,m}\|f\|_{1}\ \ \forall f\in L_{1}(\mathcal{N});$
@@
We now appeal to the weak type $(1,1)$ boundedness of $(H_{T})_{T}$ stated in Theorem 2.4 (i) that for any $\lambda>0$, there exists a projection $e\in\mathcal{N}$ such that $\mu\otimes\tau(1-e)\leq C_{1,m}\frac{\|f\|_{1}}{\lambda}$ and
$\sup_{0<P\leq\frac{m}{m+1}}\|eJ_{P}fe\|_{\infty}\leq\int_{0}^{\infty}\sup_{T>0}\|eH_{T}fe\|_{\infty}d\nu_{P}(T)\leq\lambda.$
@@
By the weak $(1,1)$-estimate of $T_{K}^{M}$, for any $\lambda>0$ there exists a projection $e\in\mathcal{N}$ such that
$\sup_{0\leq K\leq\frac{d}{m+1}}\|e(T_{K}^{D}f)e\|_{\infty}\leq C_{m}\sup_{L\leq\frac{md}{m+1}}\|e(T^{M}_{L}(\sigma_{d}\ast f))e\|_{\infty}\leq C_{m}\lambda$
@@
$\mu\otimes\tau(1-e)\leq\frac{\|\sigma_{d}\ast f\|_{1}}{\lambda}\leq C_{m}\frac{\|f\|_{1}}{\lambda},$
@@
The following elementary properties on $A_{n}^{\alpha}$ are well-known, see [22, Lemma 3].
@@
Let $(\mathcal{M},\tau)$ be a semifinite von Neumann algebra. For any $\alpha\in{\mathbb{C}}$ and $p>1$, there exists a constant $C_{m,p,\alpha}$ depends only on $m,p,\alpha$ such that $f\in L_{p}(\mathcal{N})=L_{p}(\mathbb{Z}_{m+1}^{d};L_{p}(\mathcal{M}))$
@@
Indeed, Theorem 4.2 directly yields estimates (12) and (13) through the identifications $M_{n}^{-1}f=T_{n}f$ and $N_{n}^{-1}f=T_{d-n}f$. Our proof strategy for Theorem 4.2 follows the framework established by Nevo and Stein in [22, Section 3], which we adapt to the noncommutative setting. The proof of Theorem 4.2 will be divided into the following four lemmas.
@@
By the convexity of the operator-valued function $x\mapsto|x|^{2}$ (see e.g. [20, (1.13)]),
@@
On the other hand, it follows from [22, Lemma 2] that $\Delta S_{k}^{-t}=S_{k}^{-t-1}$. Therefore, we find
@@
We prove the first statement, as the other one can be done similarly. Assume that $f$ is positive and $n\leq 10t$ first. By the convolution formula stated in [22, Lemma 2], one has
@@
It remains to estimate the first term. Observe that for every nonnegative integer $\ell$ (see [18, Lemma 2.4]),
@@
Let $({\mathcal{M}},\tau)$ be a semi-finite von Neumann algebra and $\alpha:{\mathbb{Z}}_{m+1}^{d}\to\text{Aut}(\mathcal{M},\tau)$ be a $\tau$-preserving action. Then for every $s\in{\mathbb{Z}}^{d}_{m+1}$, the map $\alpha_{s}$ extends to a positive isometry on $L_{p}({\mathcal{M}})$ for all $p\geq 1$. It follows from [12, Proposition 5.3] that $(\alpha_{s}\otimes id_{\ell_{\infty}})_{s\in{\mathbb{Z}}^{d}_{m+1}}$ extends to a family of isometry on $L_{p}({\mathcal{M}};\ell_{\infty})$ for $1\leq p\leq\infty$, still denoted by $(\alpha_{s})_{s}$. Hence, for each $s\in{\mathbb{Z}}^{d}_{m+1}$,
@@
Given $x$, define an operator-valued function $f\in L_{p}({\mathbb{Z}}^{d}_{m+1},L_{p}(\mathcal{M}))$ as
@@
By the definition of $\bigl\|(T_{k}f)_{1\leq k\leq d}\bigr\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}};\ell_{\infty})}$, there exists a factorization $T_{k}f=ay_{k}b$ with $a,b\in L_{2p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}})$ and $y_{k}\in L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}}$ such that
$\|a\|_{2p}\sup_{1\leq k\leq d}\|y_{k}\|_{\infty}\|b\|_{2p}\leq\bigl\|(T_{k}{f})_{1\leq k\leq d}\bigr\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}};\ell_{\infty})}+\delta.$
@@
$\leq$ $\big(\|(T_{k}{f})_{1\leq k\leq d}\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}};\ell_{\infty})}+\delta\big)^{p}.$
@@
(19) $\sum_{s\in{\mathbb{Z}}_{m+1}^{d}}\|(T_{k}{f}(s))_{1\leq k\leq d}\|_{L_{p}({\mathcal{M}};\ell_{\infty})}^{p}\leq\|(T_{k}{f})_{1\leq k\leq d}\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}};\ell_{\infty})}^{p}.$
@@
$\leq\frac{1}{|{\mathbb{Z}}^{d}_{m+1}|}\|(T_{k}f)_{1\leq k\leq d}\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\bar{\otimes}{\mathcal{M}};\ell_{\infty})}^{p}$
@@
$=C_{m,p}\frac{1}{|{\mathbb{Z}}^{d}_{m+1}|}\sum_{s\in{\mathbb{Z}}_{m+1}^{d}}\|\alpha_{s}x\|_{L_{p}({\mathcal{M}})}^{p}.$
$=C_{m,p}\|x\|_{L_{p}({\mathcal{M}})}^{p}.$
where the second inequality above uses the strong type $(p,p)$ estimate of $(T_{k})_{1\leq k\leq d}$ by Theorem 1.1, and the last equality uses the fact that $\alpha_{s}$ is isometric on $L_{p}({\mathcal{M}})$ for each $s\in{\mathbb{Z}}_{m+1}^{d}$. This proves the first estimate.
For the second assertions, let $x\in L_{p}({\mathcal{M}})$. Note that for each $k$, $M_{k}$ is unital and completely positive. Then by Kadison-Schwarz inequality for unital $2$-positive map, we have $M_{k}(x^{*}x)\leq M_{k}(x)^{*}M_{k}(x).$ Now applying the first assertion to $x^{*}x$, we deduce from Remark 2.1 that there exists positive $b\in L_{p/2}({\mathcal{M}})$ such that
@@
Hence, $M_{k}(x)^{*}M_{k}(x)\leq M_{k}(x^{*}x)\leq b$, which implies that for each $k$, there is a $y_{k}\in{\mathcal{M}}$ such that $M_{k}(x)=y_{k}b^{1/2}$ and $\|y_{k}\|_{\infty}$. This gives the desired factorization of $(M_{k}(x))_{k}$ as an element in $L_{p}({\mathcal{M}};\ell^{c}_{\infty})$ and the proof is finished. ∎

### v2

# Dimension-free maximal inequalities for noncommutative spherical means over cyclic groups

@@ Abstract.
In this paper, we establish dimension-free $L_{p}$-estimates for operator-valued maximal spherical means over cyclic groups ${\mathbb{Z}}_{m+1}^{d}$ for all $p>1$ and $m\geq 1$. The key ingredient is a noncommutative extension of the spectral technique developed by Nevo and Stein. As an application, we obtain a noncommutative spherical maximal inequality for automorphism actions on von Neumann algebras, along with several concrete examples.
@@ 1. Introduction
The purpose of this paper is to establish the dimension-free $L_{p}$-bound (1) to the operator-valued setting, which then implies the corresponding noncommutative maximal inequalities. Let $\mathcal{M}$ be a von Neumann algebra equipped with a normal semifinite faithful trace $\tau$. The noncommutative $L_{p}$-spaces associated to the pair $({\mathcal{M}},\tau)$ is denoted by $L_{p}({\mathcal{M}})$. For each integer $k$, define the convolution operator
@@
for operator-valued functions $f:{\mathbb{Z}}^{d}_{m+1}\rightarrow\mathcal{M}$. Let $\mathcal{N}=L_{\infty}({\mathbb{Z}}^{d}_{m+1},{\mathcal{M}})\cong L_{\infty}({\mathbb{Z}}^{d}_{m+1})\overline{\otimes}{\mathcal{M}}$ be the von Neumann algebra tensor product equipped with the tensor trace $\varphi=\mu\otimes\tau$, where $\mu$ is the Haar measure on ${\mathbb{Z}}^{d}_{m+1}$. Note that for $0<p<\infty$, $L_{p}(\mathcal{N})$ can be identified as the Bochner $L_{p}$-space $L_{p}({\mathbb{Z}}^{d}_{m+1};L_{p}({\mathcal{M}}))$. The following is our first main result and we refer to the next section for the precise definition of $L_{p}({\mathcal{M}};\ell_{\infty})$.
@@
Let $(T_{k})_{1\leq k\leq d}$ be the convolution operators defined as (2). Then for any $p>1$ and any operator-valued functions $f\in L_{p}(\mathcal{N})$,
@@
As noted in [14, Remark 1.6], the region $p>1$ is optimal in the sense that the optimal constant in the weak $(1,1)$ bound grows at least as fast as $\sqrt{d}$:
@@
An immediate difficulty in proving Theorem 1.1 is the fact that the classical maximal function $Mf=\sup_{1\leq k\leq d}|T_{k}f|$ does not make sense in general when $(T_{k}f)$ is a sequence of operators (see [18]). Indeed, the right formulation of noncommutative $L_{p}$-maximal inequalities was achieved via the introduction of vector-valued noncommutative $L_{p}$-spaces $L_{p}({\mathcal{M}};\ell_{\infty})$, which play the roles of the $L_{p}$-norm of maximal function $Mf$ in the noncommutative setting [17].
As an application, we establish the noncommutative spherical maximal inequality for automorphism actions of ${\mathbb{Z}}_{m+1}^{d}$ on von Neumann algebras. Recall that for a von Neumann algebra $\mathcal{M}$}, we say that $\alpha:{\mathbb{Z}}_{m+1}^{d}\rightarrow\text{Aut}(\mathcal{M})$ is an action if for each $u$, $\alpha_{u}:\mathcal{M}\to\mathcal{M}$ is a *-preserving automorphism; and for all $s,u\in{\mathbb{Z}}_{m+1}^{d}$, $\alpha_{s}\circ\alpha_{u}=\alpha_{s+u}$. If in additional, $\tau\circ\alpha_{u}=\tau$ for all $u\in{\mathbb{Z}}_{m+1}^{d}$, we say $\alpha$ is a $\tau$-preserving action or an action of $\tau$-preserving automorphisms, denoted as $\alpha\curvearrowright(\mathcal{M},\tau)$. For each $1\leq k\leq d$, define
@@
Theorem 1.1, together with the noncommutative Calderón transference principle developed in [12], implies the following result. Readers may refer to Section 2 for the definition of $\ell^{c}_{\infty}$-valued $L_{p}$-space $L_{p}({\mathcal{M}};\ell^{c}_{\infty})$.
@@
Let $\alpha$ be an $\tau$-preserving action of ${\mathbb{Z}}_{m+1}^{d}$ on a semifinite von Neumann algebra $({\mathcal{M}},\tau)$. Let $(M_{k})_{1\leq k\leq d}$ be the defined as (3). Then for $p>1$,
@@
$\|(M_{k}x)_{1\leq k\leq d}\|_{L_{p}({\mathcal{M}};\ell^{c}_{\infty})}\leq\sqrt{C_{p,m}}\|x\|_{L_{p}({\mathcal{M}})},\ \forall x\in L_{p}({\mathcal{M}}),$
@@
Theorem 1.3 can be rephrased as a maximal inequality for multiple ergodic averages. Let $(L_{i})_{1\leq i\leq d}$ be a commuting family (i.e. $L_{i}L_{j}=L_{j}L_{i}$) of $\tau$-preserving automorphisms on $\mathcal{M}$} such that $L_{i}^{m+1}=\text{id}$. For each positive integer $k$, define
@@
We conclude this introduction by outlining the structure of the paper. Section 2 reviews some basic background on noncommutative $L_{p}$-spaces. The proof of Theorem 1.1 are presented in Sections 3 and 4. Theorem 1.3 is proved in Section 5, and several illustrative examples of its applications are provided in Section 6.
@@ 2.1. Noncommutative $L_{p}$-spaces
Let $\mathcal{M}$} be a von Neumann algebra equipped with a normal semifinite faithful trace $\tau:{\mathcal{M}}_{+}\rightarrow[0,\infty]$. Let $\mathcal{S_{{\mathcal{M}}+}}$ be the set of all positive elements $x\in{\mathcal{M}}$ such that $\tau(\mathrm{supp}(x))<\infty$, where $\mathrm{supp}(x)$ denotes the support projection of $x$ and $\mathcal{S}_{{\mathcal{M}}}$ be the linear span of $\mathcal{S_{{\mathcal{M}}+}}$. Then $\mathcal{S}_{{\mathcal{M}}}$ forms a $w^{*}$-dense $\ast$-subalgebra of $\mathcal{M}$}. For $1\leq p<\infty$, the noncommutative $L_{p}$-space $L_{p}({\mathcal{M}},\tau)$ (in short, $L_{p}(\mathcal{M})$) is the completion of $\mathcal{S}_{{\mathcal{M}}}$ under the norm,
@@ 2.2. Vector-valued noncommutative $L_{p}$-spaces
    $L_{p}({\mathcal{M}};\ell_{1})=\big(L_{p_{0}}({\mathcal{M}};\ell_{1}),L_{p_{1}}({\mathcal{M}};\ell_{1})\big)_{\theta}\ \ \text{isometrically},$
@@
Finally, we introduce the $\ell_{\infty}$-valued noncommutative weak $L_{p}$-spaces $\mathcal{L}_{p,\infty}({\mathcal{M}};\ell_{\infty})$ (see e.g. [11]). Given a sequence $(x_{n})_{n}$ in $L_{p}(\mathcal{M})$ with $1\leq p<\infty$, we define
$\|(x_{n})_{n}\|_{\mathcal{L}_{p,\infty}({\mathcal{M}};\ell_{\infty})}=\sup_{\mathcal{L}>0}\ \mathcal{L}\inf\Big\{\big(\tau(e^{\perp})\big)^{\frac{1}{p}}:e\in\mathcal{P}({\mathcal{M}})\text{ s.t. }\|ex_{k}e\|_{\infty}\leq\mathcal{L}\ \mbox{for\ all}\ k\Big\},$
where $\mathcal{P}({\mathcal{M}})$ is the set of all projections in $\mathcal{M}$}. The space $\mathcal{L}_{p,\infty}({\mathcal{M}};\ell_{\infty})$ consists of all sequences $x=(x_{n})_{n}$ for which this quasi-norm is finite.
@@ 2.3. Noise operators
    $\|(H_{T}f)_{T>0}\|_{\mathcal{L}_{1,\infty}(\mathcal{N};\ell_{\infty})}\leq C_{1,m}\|f\|_{L_{1}(\mathcal{N})};$
@@ 3. The smooth spherical maximal operators
    $\|(T_{K}^{M}f)_{0\leq K\leq\frac{md}{m+1}}\|_{\mathcal{L}_{1,\infty}(\mathcal{N};\ell_{\infty})}\leq C_{1,m}\|f\|_{L_{1}(\mathcal{N})}\ ,$
    $\|(T_{K}^{D}f)_{0\leq K\leq\frac{d}{m+1}}\|_{\mathcal{L}_{1,\infty}(\mathcal{N};\ell_{\infty})}\leq C_{1,m}\|f\|_{L_{1}(\mathcal{N})}\ ;$
@@
    $\|(J_{P}f)_{0<P\leq\frac{m}{m+1}}\|_{\mathcal{L}_{1,\infty}(\mathcal{N};\ell_{\infty})}\leq C_{1,m}\|f\|_{1}\ \ \forall f\in L_{1}(\mathcal{N});$
@@
We now appeal to the weak type $(1,1)$ boundedness of $(H_{T})_{T}$ stated in Theorem 2.4 (i) that for any $\mathcal{L}>0$, there exists a projection $e\in\mathcal{N}$ such that $\mu\otimes\tau(1-e)\leq C_{1,m}\frac{\|f\|_{1}}{\mathcal{L}}$ and
$\sup_{0<P\leq\frac{m}{m+1}}\|eJ_{P}fe\|_{\infty}\leq\int_{0}^{\infty}\sup_{T>0}\|eH_{T}fe\|_{\infty}d\nu_{P}(T)\leq\mathcal{L}.$
@@
By the weak $(1,1)$-estimate of $T_{K}^{M}$, for any $\mathcal{L}>0$ there exists a projection $e\in\mathcal{N}$ such that
$\sup_{0\leq K\leq\frac{d}{m+1}}\|e(T_{K}^{D}f)e\|_{\infty}\leq C_{m}\sup_{L\leq\frac{md}{m+1}}\|e(T^{M}_{L}(\sigma_{d}\ast f))e\|_{\infty}\leq C_{m}\mathcal{L}$
@@
$\mu\otimes\tau(1-e)\leq\frac{\|\sigma_{d}\ast f\|_{1}}{\mathcal{L}}\leq C_{m}\frac{\|f\|_{1}}{\mathcal{L}},$
@@
The following elementary properties on $A_{n}^{\alpha}$ are well-known, see [23, Lemma 3].
@@
Let $(\mathcal{M},\tau)$ be a semifinite von Neumann algebra. For any $\alpha\in{\mathbb{C}}$ and $p>1$, there exists a constant $C_{m,p,\alpha}$ depends only on $m,p,\alpha$ such that $f\in L_{p}(\mathcal{N})$
@@
Indeed, Theorem 4.2 directly yields estimates (12) and (13) through the identifications $M_{n}^{-1}f=T_{n}f$ and $N_{n}^{-1}f=T_{d-n}f$. Our proof strategy for Theorem 4.2 follows the framework established by Nevo and Stein in [23, Section 3], which we adapt to the noncommutative setting. The proof of Theorem 4.2 will be divided into the following four lemmas.
@@
By the convexity of the operator-valued function $x\mapsto|x|^{2}$ (see e.g. [21, (1.13)]),
@@
On the other hand, it follows from [23, Lemma 2] that $\Delta S_{k}^{-t}=S_{k}^{-t-1}$. Therefore, we find
@@
We prove the first statement, as the other one can be done similarly. Assume that $f$ is positive and $n\leq 10t$ first. By the convolution formula stated in [23, Lemma 2], one has
@@
It remains to estimate the first term. Observe that for every nonnegative integer $\ell$ (see [19, Lemma 2.4]),
@@
Let $({\mathcal{M}},\tau)$ be a semifinite von Neumann algebra and $\alpha:{\mathbb{Z}}_{m+1}^{d}\to\text{Aut}(\mathcal{M},\tau)$ be a $\tau$-preserving action. Then for every $s\in{\mathbb{Z}}^{d}_{m+1}$, the map $\alpha_{s}$ extends to a positive isometry on $L_{p}({\mathcal{M}})$ for all $p\geq 1$. It follows from [13, Proposition 5.3] that $(\alpha_{s}\otimes id_{\ell_{\infty}})_{s\in{\mathbb{Z}}^{d}_{m+1}}$ extends to a family of isometry on $L_{p}({\mathcal{M}};\ell_{\infty})$ for $1\leq p\leq\infty$, still denoted by $(\alpha_{s})_{s}$. Hence, for each $s\in{\mathbb{Z}}^{d}_{m+1}$,
@@
Given $x$, define an operator-valued function $f\in L_{p}({\mathbb{Z}}^{d}_{m+1};L_{p}(\mathcal{M}))$ as
@@
By the definition of $\bigl\|(T_{k}f)_{1\leq k\leq d}\bigr\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\overline{\otimes}{\mathcal{M}};\ell_{\infty})}$, there exists a factorization $T_{k}f=ay_{k}b$ with $a,b\in L_{2p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\overline{\otimes}{\mathcal{M}})$ and $y_{k}\in L_{\infty}({\mathbb{Z}}_{m+1}^{d})\overline{\otimes}{\mathcal{M}}$ such that
$\|a\|_{2p}\sup_{1\leq k\leq d}\|y_{k}\|_{\infty}\|b\|_{2p}\leq\bigl\|(T_{k}{f})_{1\leq k\leq d}\bigr\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\overline{\otimes}{\mathcal{M}};\ell_{\infty})}+\delta.$
@@
$\leq$ $\big(\|(T_{k}{f})_{1\leq k\leq d}\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\overline{\otimes}{\mathcal{M}};\ell_{\infty})}+\delta\big)^{p}.$
@@
(19) $\sum_{s\in{\mathbb{Z}}_{m+1}^{d}}\|(T_{k}{f}(s))_{1\leq k\leq d}\|_{L_{p}({\mathcal{M}};\ell_{\infty})}^{p}\leq\|(T_{k}{f})_{1\leq k\leq d}\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\overline{\otimes}{\mathcal{M}};\ell_{\infty})}^{p}.$
@@
$\leq\frac{1}{|{\mathbb{Z}}^{d}_{m+1}|}\|(T_{k}f)_{1\leq k\leq d}\|_{L_{p}(L_{\infty}({\mathbb{Z}}_{m+1}^{d})\overline{\otimes}{\mathcal{M}};\ell_{\infty})}^{p}$
@@
$=C_{m,p}\frac{1}{|{\mathbb{Z}}^{d}_{m+1}|}\sum_{s\in{\mathbb{Z}}_{m+1}^{d}}\|\alpha_{s}x\|_{L_{p}({\mathcal{M}})}^{p}$
$=C_{m,p}\|x\|_{L_{p}({\mathcal{M}})}^{p},$
where the second inequality above uses the strong type $(p,p)$ estimate of $(T_{k})_{1\leq k\leq d}$ by Theorem 1.1, and the last equality follows from the fact that $\alpha_{s}$ is isometric on $L_{p}({\mathcal{M}})$ for each $s\in{\mathbb{Z}}_{m+1}^{d}$. This proves the first estimate.
For the second assertion, let $x\in L_{p}({\mathcal{M}})$. Note that for each $k$, $M_{k}$ is unital and completely positive. Then by Kadison-Schwarz inequality for unital $2$-positive map, we have $M_{k}(x^{*}x)\leq M_{k}(x)^{*}M_{k}(x).$ Now applying the first assertion to $x^{*}x$, we deduce from Remark 2.1 that there exists positive $b\in L_{p/2}({\mathcal{M}})$ such that
@@
Hence, $M_{k}(x)^{*}M_{k}(x)\leq M_{k}(x^{*}x)\leq b$, which implies that for each $k$, there is a $y_{k}\in{\mathcal{M}}$ such that $M_{k}(x)=y_{k}b^{1/2}$ for all $k$. This gives the desired factorization of $(M_{k}(x))_{k}$ as an element in $L_{p}({\mathcal{M}};\ell^{c}_{\infty})$ and the proof is finished. ∎